# A-Bench: Are LMMs Masters at Evaluating AI-generated Images?

### v1

@@ Abstract
How to accurately and efficiently assess AI-generated images (AIGIs) remains a critical challenge for generative models. Given the high costs and extensive time commitments required for user studies, many researchers have turned towards employing large multi-modal models (LMMs) as AIGI evaluators, the precision and validity of which are still questionable. Furthermore, traditional benchmarks often utilize mostly natural-captured content rather than AIGIs to test the abilities of LMMs, leading to a noticeable gap for AIGIs. Therefore, we introduce A-Bench in this paper, a benchmark designed to diagnose whether LMMs are masters at evaluating AIGIs. Specifically, A-Bench is organized under two key principles: 1) Emphasizing both high-level semantic understanding and low-level visual quality perception to address the intricate demands of AIGIs. 2) Various generative models are utilized for AIGI creation, and various LMMs are employed for evaluation, which ensures a comprehensive validation scope. Ultimately, 2,864 AIGIs from 16 text-to-image models are sampled, each paired with question-answers annotated by human experts, and tested across 18 leading LMMs. We hope that A-Bench will significantly enhance the evaluation process and promote the generation quality for AIGIs. The benchmark is available at https://github.com/Q-Future/A-Bench.
@@ 1 Introduction
One look is worth a thousand words. Inspired by this age-old adage, numerous researchers dedicate their efforts to developing text-to-image (T2I) models that vividly bring text to life through imagery. These T2I models, driven by free-form text prompts, aim to create images that accurately align with the text and showcase high perceptual quality. Innovations such as AlignDRAW (Mansimov et al., 2015) and the text-conditional GAN (Reed et al., 2016) have introduced differential architecture for image generation. The field continues to advance with the development of stable diffusion models (Saharia et al., 2022; Rombach et al., 2022a), significantly propelling T2I technology forward. On the commercial front, major corporations leverage vast-scale data to launch stunningly effective T2I models, such as DALL-E (Ramesh et al., 2022), Midjourney (Holz, 2023), Parti (Yu et al., 2022), etc. However, despite their diversity and widespread adoption, all these advanced T2I models occasionally face issues of low alignment with prompts and low perceptual quality in creating AI-generated images (AIGIs), necessitating careful evaluation and improvement.
The alignment and quality evaluation of AIGIs present significant challenges that small expert models attempt to address. Although these small expert models offer some solutions, they possess inherent drawbacks and often fail to meet contemporary demands. Specifically, for alignment assessment, CLIP-based similarity models struggle with accurately judging alignment, particularly with complex text prompts (Radford et al., 2021a). When it comes to quality evaluation, traditional image quality/aesthetic assessment methods (IQA/IAA) are not capable of identifying AIGI-generative distortions (Wu et al., 2023a, b), rendering them unsuitable for this specialized task.
@@
Many researchers are increasingly relying on large language models (LLMs) and large multi-modal models (LMMs) for their human-like processing capabilities, which are presumed to enable accurate judgments of alignment and quality in generated content. Consequently, a variety of LMM-based evaluators have been developed, including VIE-Score (Ku et al., 2023), Prometheus (Kim et al., 2023), GPT4V-Eval (Lin et al., 2024), TIFA (Hu et al., 2023), and Davidsonian Graph (Cho et al., 2023), among others. Despite these advancements, a fundamental question remains:
@@
These LMM-based metrics traditionally employ evaluation criteria such as SRCC/PLCC to determine their reliability. However, this approach only reveals how well the metrics perform, without shedding light on their specific strengths and weaknesses. To address this gap, we propose conducting a detailed and comprehensive ‘diagnostic’ benchmark $\rightarrow$ A-Bench, focusing on LMMs’ capabilities in semantic understanding and quality assessment. Rather than directly evaluating these LMM-based metrics, we focus on studying the LMMs themselves behind. We move away from computing SRCC/PLCC criteria and instead examine whether the fundamental perceptual questions can be correctly answered, which is the core basis of all LMM-based evaluators. To initiate our exploration on the AIGI evaluation abilities of LMMs, we first construct the A-Bench centered on a pivotal question:
@@ 2.1 Large Muli-modal Models
Large language models (LLMs), such as GPT-4 (OpenAI, 2023), T5 (Chung et al., 2022), and LLaMA (Touvron et al., 2023), exhibit exceptional linguistic capabilities in general human knowledge domains. By integrating visual input via CLIP (Radford et al., 2021b) and additional adaptation modules, large multi-modal models (LMMs) (Li et al., 2023a; Gao et al., 2023; Liu et al., 2023a; Dai et al., 2023; Zhang et al., 2023) are capable of addressing diverse multi-modal tasks, including image captioning, visual question answering, visual segmentation, visual classification, visual reasoning, etc. Namely, OpenFlamingo (Awadalla et al., 2023) initially integrates several gated cross-attention dense blocks into the pretrained language encoder layers. InstructBLIP (Dai et al., 2023) extends BLIP-2 (Li et al., 2023b) by incorporating vision-language instruction tuning. To further develop open-source LMMs, many works have employed GPT-4 (OpenAI, 2023) to create data for vision-language tuning, such as LLaVA series (Liu et al., 2023a, b, 2024). However, whether these LMMs are masters at evaluating T2I AIGIs is still questionable, which needs further investigation.
@@ 2.2 Multi-modal Benchmarks
Benchmarks such as COCO Caption (Chen et al., 2015) and Nocaps (Agrawal et al., 2019) evaluate the capability of models to generate textual descriptions for images. Subsequently, benchmarks like GQA (Hudson and Manning, 2019) and OK-VQA (Marino et al., 2019) focus on visual question answering, assessing multi-modal models’ visual perception and reasoning abilities. Further complexities are added in benchmarks such as TextVQA (Singh et al., 2019) and ScienceQA (Lu et al., 2022), which incorporate OCR tasks and commonsense reasoning, respectively. MME (Fu et al., 2023) and MMbench (Liu et al., 2023c) provide comprehensive evaluations of LMMs across various subtasks. Additionally, MMMU (Yue et al., 2023) targets extensive multi-disciplinary tasks that require college-level knowledge and sophisticated reasoning. More recently, Q-Bench (Wu et al., 2023a) focuses specifically on assessing the low-level visual perception capabilities of LMMs. Despite these efforts, there is still a gap in systematic benchmarks for assessing the abilities of LMMs in AIGI evaluation, prompting the development of A-Bench to address this shortfall.
@@ 3.2.1 High-level Semantic Understanding
This aspect concentrates on the fundamental semantic understanding of AIGIs, which can be subdivided into two distinct areas: 1) Major Object Recognition, which involves recognizing the primary objects in the image, such as humans or objects depicted in the foreground. 2) Minor Object Recognition, which pertains to the identification of less-prominent objects within the image, such as background elements or secondary characters.
@@
This dimension focuses on the discriminative semantic understanding of AIGIs crafted with Bag-of-words prompts (encompassing rich descriptive attributes or complex object relationships (Qu et al., 2024)). This can be subdivided into the following aspects: 1) Attributes Awareness, defined as the capability to accurately identify the attributes of objects in AIGIs. 2) Additionally, given that T2I models may incorrectly interpret nouns as adjectives, resulting in the unwanted generation of objects instead of the intended attributes, we have also introduced a dimension called Nouns as Adjectives Awareness to address this issue. 3) Composition Identification, recognized as the ability to correctly comprehend the compositional relationships such as orientation, occlusion, size comparison, and spatial arrangement. 4) Number of Objects Counting, regarded as the ability to accurately count the specified objects in the image, which is crucial for assessing whether the AIGI aligns with the numerical specifications of the prompt.
@@
This aspect emphasizes the reasoning ability to utilize external knowledge not directly depicted in the images (Schwenk et al., 2022), and can be broken down into the following dimensions: 1) Specific Terms Recognition: This involves identifying specific scenes and objects related to distinct domains such as geography, sports, science, materials, food, everyday life, creatures, brands, and styles. 2) Contradiction Overcome, recognized as the ability to correctly interpret AIGIs even when their content contradicts established world knowledge, which is particularly crucial for evaluating AIGIs generated from controversial prompts.
@@ 3.2.2 Low-level Quality Perception
Conversely, to determine the ability of LMMs on image visual quality, we conduct the A-BenchP2, comprising 1,456 challenging multi-modal question-answer pairs centered on low-level quality perception for AIGIs, which can be categorized into the following aspects: 1) Technical Quality Perception This indicates the low-level characteristics that directly degrade the quality of images, such as blur, noise, exposure, etc (Su et al., 2021; Ying et al., 2020). 2) Aesthetic Quality Evaluation This indicates the attributes that affect the aesthetic appeal of AIGIs and evoke varied human feelings, which include color, lighting, etc (Huang et al., 2024). 3) Generative Distortion Assessment This indicates the unexpected AIGI-specific distortions (Chen et al., 2023a; Li et al., 2023c, 2024), such as generative blur caused by low completion, confusing geometry structure, unnaturalness, etc.
@@ 3.3 Question Collection
#User: What painting style is represented in the image? <|IMAGE_TOKEN|>
@@ 4.1 Benchmark Candidates
To ensure the results are comprehensive and up-to-date, we select the latest and widely used LMMs for benchmarking. The Proprietary LMMs (closed-source) include Gemini 1.5 Pro (Reid et al., 2024), GPT-4v (OpenAI, 2023), GPT-4o (OpenAI, 2024), and Qwen-VL-Max (Bai et al., 2023). The Open-source LMMs include CogVLM2-19B (Llama3-8B) (Wang et al., 2023), IDEFICS-2 (Mistral-7B-Instruct-v0.2) (Huggingface, 2023), DeepSeek-VL-7B (Lu et al., 2024), InternLM-XComposer2-VL (Dong et al., 2024), LLaVA-NeXT (Llama3-8B), LLaVA-NeXT (Qwen-72B), LLaVA-NeXT (Qwen-110B) (Liu et al., 2024), mPLUG-Owl2 (LLaMA-7B) (Ye et al., 2023), LLaVA-v1.5 (Vicuna-v1.5-7B), LLaVA-v1.5 (Vicuna-v1.5-13B) (Liu et al., 2023b), CogVLM-17B (Vicuna-v1.5-7B) (Wang et al., 2023), Qwen-VL (Qwen-7B) (Bai et al., 2023), BakLLava (Mistral-7B) (Liu et al., 2023a), and Fuyu-8B (Persimmon-8B) (Adept, 2023). All LMMs are tested with zero-shot setting.
For human performance on A-Bench, we conduct a user-study experiment with five ordinary people in a controlled laboratory setting. Initially, participants familiarize themselves with the tasks through exposure to similar cases. Subsequently, they select the appropriate responses for the questions posed in the A-Bench. To maintain consistency with the conditions experienced by LMMs, the order of questions is randomized, and participants receive no additional information beyond the AIGIs, questions, and answer options. The best and worst performance is included for comparison.
@@
Figure 5: A Quick Look of the A-Bench outcomes. (a) showcases a comparative analysis of the overall accuracy between humans, 18 selected LMMs (both closed-source and open-source), and random guess. (b) displays a radar chart that details the accuracy performance (subtracting the accuracy of random guess) of the top-7 LMMs across various subcategories of A-Bench.
@@ 4.2 Findings of A-Bench
General Observation: Human > Proprietary LMMs > Open-source LMMs
A concise overview of the A-Bench results is provided in Fig. 5, revealing several general insights: 1) All LMMs significantly outperform the random guess, indicating their capabilities in handling AIGI evaluation, with Gemini 1.5 Pro leading, closely followed by GPT-4o and Qwen-VL-Max. Notably, among the open-source LMMs, which are preferred for AIGI evaluations due to their accessibility and modifiability, LLaVA-NeXT (Qwen-110B) stands out, though it still significantly lags behind closed-source competitors. 2) Even the lowest performance by humans surpasses that of all LMMs, with a noticeable 14.70% gap compared to the top-performing LMM, Gemini 1.5 Pro, indicating that LMMs are still far from adequately performing AIGI evaluation as humans. 3) A closer examination of the radar chart in Fig. 5 (b) shows that top LMMs exhibit varied performances across different sub-categories, suggesting a lack of robustness, while humans show more consistent and balanced performance across these categories, highlighting areas where LMMs need further improvement.
@@
Human (Worst) 95.20% 94.27% 96.83% 88.64% 85.54% 82.50% 81.79% 88.89% 92.39%
Human (Best) 95.42% 95.18% 99.46% 95.12% 93.42% 91.67% 84.23% 96.00% 94.00%
@@
IDEFICS-2 (Mistral-7B-Instruct-v0.2) 89.95% 91.94% 86.43% 75.51% 61.29% 71.11% 73.22% 62.86% 80.00%
DeepSeek-VL-7B 91.49% 91.13% 82.41% 83.67% 63.44% 70.00% 75.41% 60.00% 81.24%
InternLM-XComposer2-VL (InternLM2) 92.78% 95.16% 86.43% 82.65% 68.82% 72.22% 70.77% 64.29% 81.90%
LLaVA-NeXT (Llama3-8B) 92.75% 92.37% 91.12% 83.65% 61.00% 67.02% 76.20% 62.97% 82.91%
LLaVA-NeXT (Qwen-72B) 94.33% 92.74% 91.46% 81.63% 62.37% 73.33% 77.05% 61.43% 83.98%
LLaVA-NeXT (Qwen-110B) 93.81% 91.13% 90.45% 84.69% 67.74% 67.78% 76.23% 64.29% 83.62%
mPLUG-Owl2 (LLaMA-7B) 85.31% 86.29% 83.92% 79.59% 53.76% 57.78% 71.04% 58.57% 76.44%
LLaVA-v1.5 (Vicuna-v1.5-7B) 87.89% 88.71% 83.92% 75.51% 61.29% 65.56% 74.86% 62.86% 79.16%
LLaVA-v1.5 (Vicuna-v1.5-13B) 88.40% 89.52% 86.43% 79.59% 62.37% 58.89% 74.86% 61.43% 79.67%
CogVLM-17B (Vicuna-v1.5-7B) 90.46% 95.14% 85.93% 77.55% 49.46% 47.78% 73.22% 61.43% 78.52%
Qwen-VL (Qwen-7B) 86.34% 86.29% 81.41% 77.55% 52.69% 61.11% 71.58% 57.14% 76.43%
BakLLava (Mistral-7B) 88.97% 81.29% 77.36% 73.87% 52.21% 62.38% 68.49% 49.12% 74.27%
Fuyu-8B (Persimmon-8B) 81.24% 68.07% 66.67% 57.34% 42.14% 48.46% 61.06% 29.79% 63.05%
random guess 32.28% 37.24% 31.00% 42.79% 29.81% 29.80% 26.46% 32.11% 30.87%
@@
The performance results of LMMs on the A-BenchP1 subset, as detailed in Table 1, reveal several key insights: 1) Almost all LMMs show good performance in Basic Recognition, suggesting that they are quite adept at fundamental semantic understanding, which includes recognizing foreground and background objects in AIGIs. 2) However, their effectiveness diminishes in more complex tasks such as Bag-of-Words, particularly in subcategories like Nouns as Adjectives Awareness, Composition Identification, and Number of Objects Counting. These areas require deeper semantic understanding and reasoning, which is critical as users often employ complex prompts that include such nuanced elements. The LMMs’ underperformance here indicates potential challenges in accurately aligning AIGIs with user prompts. 3) Additionally, Outside Knowledge poses significant challenges, with LMMs generally achieving unsatisfactory performance in the Contradiction Overcome subcategory, where AIGIs contain content that defies common sense, requiring LMMs to override their prior knowledge to respond correctly. The subcategory Specific Terms tests the knowledge base of LMMs, where proprietary LMMs generally perform better due to being trained on more recent and extensive datasets.
@@
The performance results of LMMs on the A-BenchP2 subset, as shown in Table 2, illustrate a notable disparity in capabilities: 1) There is a significant performance gap of approximately 20% between the top-performing LMMs and human evaluators, highlighting that LMMs lag considerably in quality perception and struggle to accurately assess the quality of AIGIs. 2) Furthermore, most LMMs exhibit their weakest performance in the Generative Distortion Assessment subcategory, suggesting their ineffectiveness at identifying unexpected generative distortions, such as unnatural appearances and incorrect geometric structures. 3) Interestingly, while humans generally perform better in Technical Quality Perception compared to Aesthetic Quality Evaluation, LMMs show similar performance levels in both subcategories. This difference likely stems from the more objective nature of technical quality assessments, which leads to more consistent evaluations among humans, whereas aesthetic quality, being more subjective, results in a broader range of opinions and consequently lower performance scores.
@@
Human (Worst) 94.40% 84.41% 86.30% 90.53%
Human (Best) 94.99% 86.12% 93.04% 92.25%

### v2

@@ Abstract
How to accurately and efficiently assess AI-generated images (AIGIs) remains a critical challenge for generative models. Given the high costs and extensive time commitments of user studies, many researchers have turned towards employing large multi-modal models (LMMs) as AIGI evaluators, the precision and validity of which are still questionable. Furthermore, traditional benchmarks often utilize mostly natural-captured content rather than AIGIs to test the abilities of LMMs, leading to a noticeable gap for AIGIs. Therefore, we introduce A-Bench in this paper, a benchmark designed to diagnose whether LMMs are masters at evaluating AIGIs. Specifically, A-Bench is organized under two key principles: 1) Emphasizing both high-level semantic understanding and low-level visual quality perception to address the intricate demands of AIGIs. 2) Various generative models are utilized for AIGI creation, and various LMMs are employed for evaluation, which ensures a comprehensive validation scope. Ultimately, 2,864 AIGIs from 16 text-to-image models are sampled, each paired with question-answers annotated by human experts. We hope that A-Bench will significantly enhance the evaluation process and promote the generation quality for AIGIs.
@@ 1 Introduction
One look is worth a thousand words. Inspired by this age-old adage, numerous researchers dedicate their efforts to developing text-to-image (T2I) models that vividly bring text to life through imagery. These T2I models, driven by free-form text prompts, aim to create images that accurately align with the text and showcase high perceptual quality. Innovations such as AlignDRAW (Mansimov et al., 2015) and the text-conditional GAN (Reed et al., 2016) have introduced differential architecture for image generation. The field continues to advance with the development of stable diffusion models (Saharia et al., 2022; Rombach et al., 2022b), significantly propelling T2I technology forward. On the commercial front, major corporations leverage vast-scale data to launch stunningly effective T2I models, such as DALL-E (Ramesh et al., 2022), Midjourney (Holz, 2023), Parti (Yu et al., 2022), etc. However, despite their diversity and widespread adoption, all these advanced T2I models occasionally face issues of low alignment with prompts and low perceptual quality in creating AI-generated images (AIGIs), necessitating careful evaluation and improvement.
The alignment and quality evaluation of AIGIs present significant challenges that small expert models attempt to address. Although these small expert models offer some solutions, they possess inherent drawbacks and often fail to meet contemporary demands. Specifically, for alignment assessment, CLIP-based similarity models struggle with accurately judging alignment, particularly with complex text prompts (Radford et al., 2021b). When it comes to quality evaluation, traditional image quality/aesthetic assessment methods (IQA/IAA) are not capable of identifying AIGI-generative distortions (Wu et al., 2023a; b), rendering them unsuitable for this specialized task.
@@
Many researchers are increasingly relying on large language models (LLMs) and large multi-modal models (LMMs) for their human-like processing capabilities, which are presumed to enable accurate judgments of alignment and quality in generated content. Consequently, many LMM-based evaluators have been developed, including VIE-Score (Ku et al., 2023), Prometheus (Kim et al., 2023), VQAScore (Lin et al., 2024), GPT4V-Eval (Zhang et al., 2023b), TIFA (Hu et al., 2023), and Davidsonian Graph (Cho et al., 2023), etc. However, a fundamental question remains:
@@
These LMM-based metrics traditionally employ evaluation criteria such as SRCC/PLCC to determine their reliability. However, this approach only reveals how well the metrics perform, without shedding light on their specific strengths and weaknesses. To address this gap, we propose conducting a comprehensive ‘diagnostic’ benchmark $\rightarrow$ A-Bench, focusing on LMMs’ capabilities in semantic understanding and quality assessment. Rather than directly evaluating these LMM-based metrics, we focus on the LMMs themselves behind. We move away from computing SRCC/PLCC criteria and instead examine whether the fundamental perceptual questions can be correctly answered, which is the core basis of all LMM-based evaluators. To initiate our exploration on the AIGI evaluation abilities of LMMs, we first construct the A-Bench centered on a pivotal question:
@@ 2.1 Large Muli-modal Models
Large language models (LLMs), such as GPT-4 (OpenAI, 2023), T5 (Chung et al., 2022), and LLaMA (Touvron et al., 2023), exhibit exceptional linguistic capabilities in general human knowledge domains. By integrating visual input via CLIP (Radford et al., 2021a) and additional adaptation modules, large multi-modal models (LMMs) (Li et al., 2023a; Gao et al., 2023; Liu et al., 2023b; Dai et al., 2023; Zhang et al., 2023a) are capable of addressing diverse multi-modal tasks, including image captioning, visual question answering, visual segmentation, visual classification, visual reasoning, etc. Namely, OpenFlamingo (Awadalla et al., 2023) initially integrates several gated cross-attention dense blocks into the pretrained language encoder layers. InstructBLIP (Dai et al., 2023) extends BLIP-2 (Li et al., 2023c) by incorporating vision-language instruction tuning. To further develop open-source LMMs, many works have employed GPT-4 (OpenAI, 2023) to create data for vision-language tuning, such as LLaVA series (Liu et al., 2023b; a; 2024). However, whether these LMMs are masters at evaluating T2I AIGIs is still questionable, which needs further investigation.
@@ 2.2 Multi-modal Benchmarks
Benchmarks such as COCO Caption (Chen et al., 2015) and Nocaps (Agrawal et al., 2019) evaluate the capability of models to generate textual descriptions for images. Subsequently, benchmarks like GQA (Hudson & Manning, 2019) and OK-VQA (Marino et al., 2019) focus on visual question answering, assessing multi-modal models’ visual perception and reasoning abilities. Further complexities are added in benchmarks such as TextVQA (Singh et al., 2019) and ScienceQA (Lu et al., 2022), which incorporate OCR tasks and commonsense reasoning, respectively. MME (Fu et al., 2023) and MMbench (Liu et al., 2023c) provide comprehensive evaluations of LMMs across various subtasks. Additionally, MMMU (Yue et al., 2023) targets extensive multi-disciplinary tasks that require college-level knowledge and sophisticated reasoning. More recently, Q-Bench (Wu et al., 2023a) focuses specifically on assessing the low-level visual perception capabilities of LMMs. Despite these efforts, there is still a gap in systematic benchmarks for assessing the abilities of LMMs in AIGI evaluation, prompting the development of A-Bench to address this shortfall.
@@ 3.2.1 High-level Semantic Understanding
This aspect concentrates on the fundamental semantic understanding of AIGIs (Nichol et al., 2021; Saharia et al., 2022), which can be subdivided into two distinct areas based on the objects of interest: 1) Major Object Recognition, which involves recognizing the primary objects in the image, such as humans or objects depicted in the foreground. 2) Minor Object Recognition, which pertains to the identification of less-prominent objects within the image, such as background elements or secondary characters.
@@
This dimension focuses on the discriminative semantic understanding of AIGIs crafted with Bag-of-words prompts (encompassing rich descriptive attributes or complex object relationships (Qu et al., 2024)). This can be subdivided into the following aspects related to the crucial points of T2I generation alignment: 1) Attributes Awareness, defined as the capability to accurately identify the attributes of objects in AIGIs (Xu et al., 2023; Liu et al., 2023c). 2) Additionally, given that T2I models may incorrectly interpret nouns as adjectives, resulting in the unwanted generation of objects instead of the intended attributes (Chatterjee et al., 2025; Motamed et al., 2023), we have also introduced a dimension called Nouns as Adjectives Awareness to address this issue. 3) Composition Identification, recognized as the ability to correctly comprehend the compositional relationships such as orientation, occlusion, size comparison, and spatial arrangement (Wang et al., 2024b; Zhang et al., 2024). 4) Number of Objects Counting, regarded as the ability to accurately count the specified objects in the image, which is crucial for assessing whether the AIGI aligns with the numerical specifications of the prompt (Binyamin et al., 2024).
@@
This aspect emphasizes the reasoning ability to utilize external knowledge not directly depicted in the images (Schwenk et al., 2022), and can be broken down into the following dimensions: 1) Specific Terms Recognition: This involves identifying specific scenes and objects related to distinct domains such as geography, sports, science, materials, food, everyday life, creatures, brands, and styles. 2) Contradiction Overcome, recognized as the ability to correctly interpret AIGIs even when their content contradicts established world knowledge, which is particularly crucial for evaluating AIGIs generated from controversial prompts (Hou et al., 2024).
@@ 3.2.2 Low-level Quality Perception
Conversely, to determine the ability of LMMs on image visual quality, we conduct the A-BenchP2, comprising 1,456 challenging multi-modal question-answer pairs centered on low-level quality perception for AIGIs, which can be categorized into the following aspects: 1) Technical Quality Perception This indicates the low-level characteristics that directly degrade the quality of images, such as blur, noise, exposure, etc (Su et al., 2021; Ying et al., 2020). 2) Aesthetic Quality Evaluation This indicates the attributes that affect the aesthetic appeal of AIGIs and evoke varied human feelings, which include color, lighting, etc (Huang et al., 2024). 3) Generative Distortion Assessment This indicates the unexpected AIGI-specific distortions (Chen et al., 2023b; Li et al., 2023b; 2024), such as generative blur caused by low completion, confusing geometry structure, unnaturalness, etc.
@@ 3.3 Question Collection
#User: What painting style is represented in the image? $|IMAGE\_TOKEN|$
@@ 4.1 Benchmark Candidates
To ensure the results are comprehensive and up-to-date, we select the widely used LMMs for benchmarking. The Proprietary LMMs (closed-source) include Gemini 1.5 Pro (Reid et al., 2024), GPT-4v (OpenAI, 2023), GPT-4o (2024-05-13) (OpenAI, 2024), and Qwen-VL-Max (Bai et al., 2023). The Open-source LMMs include Qwen2-VL-72B (Qwen2-72B(Wang et al., 2024a), MiniCPM-V2.6 (Qwen2-7B(Yao et al., 2024), InternVL2-40B (Nous-Hermes-2-Yi-34B(Chen et al., 2024), Ovis1.5 (Llama3-8B), LLaVA-OneVision (Qwen2-7B(Lu et al., 2024b), CogVLM2-19B (Llama3-8B) (Wang et al., 2023), IDEFICS-2 (Mistral-7B-Instruct-v0.2) (Huggingface, 2023), DeepSeek-VL-7B (Lu et al., 2024a), InternLM-XComposer2-VL (Dong et al., 2024), LLaVA-NeXT (Llama3-8B), LLaVA-NeXT (Qwen-72B), LLaVA-NeXT (Qwen-110B) (Liu et al., 2024), mPLUG-Owl2 (LLaMA-7B) (Ye et al., 2023), LLaVA-v1.5 (Vicuna-v1.5-7B), LLaVA-v1.5 (Vicuna-v1.5-13B) (Liu et al., 2023a), CogVLM-17B (Vicuna-v1.5-7B) (Wang et al., 2023), Qwen-VL (Qwen-7B) (Bai et al., 2023), BakLLava (Mistral-7B) (Liu et al., 2023b), and Fuyu-8B (Persimmon-8B) (Adept, 2023). All LMMs are tested with zero-shot setting. It’s worth noting that the instruction prompt might slightly differ for different LMMs according to the official setting.
4.2 Human Performance
For human performance on A-Bench, we conduct a user-study experiment with five ordinary people in a controlled laboratory setting. Initially, participants familiarize themselves with the tasks through exposure to similar cases. Subsequently, they select the appropriate responses for the questions posed in the A-Bench. To maintain consistency with the conditions experienced by LMMs, the order of questions is randomized, and participants receive no additional information beyond the AIGIs, questions, and answer options. The best and worst performance is included for comparison. More details about acquiring human performance can be referred to in Sec.A.5.
@@ 4.2 Human Performance
Figure 5: A Quick Look of the A-Bench outcomes. (a) showcases a comparative analysis of the overall accuracy between humans, 23 selected LMMs (both closed-source and open-source), and random guess. (b) displays a radar chart that details the accuracy performance (subtracting the accuracy of random guess) of the top-7 LMMs across various subcategories of A-Bench.
@@ 4.3 Findings of A-Bench
General Observation: Proprietary LMMs deliver performance comparable to the best open-source LMMs.
A concise overview of the A-Bench results is provided in Fig. 5, revealing several general insights: 1) All LMMs significantly outperform the random guess, indicating their capabilities in handling AIGI evaluation, with Qwen2-VL-72B leading, closely followed by Gemini 1.5 Pro and GPT-4o. Notably, among the open-source LMMs, which are preferred for AIGI evaluations due to their accessibility and modifiability, Qwen2-VL-72B stands out, even outperforming the best closed-source competitors. 2) Even the lowest performance by humans surpasses that of all LMMs, with a noticeable 14.62% gap compared to the top-performing LMM, Qwen2-VL-72B, indicating that LMMs are still far from adequately performing AIGI evaluation as humans. 3) A closer examination of the radar chart in Fig. 5 (b) shows that top LMMs exhibit varied performances across different sub-categories, suggesting a lack of robustness, while humans show more consistent and balanced performance across these categories, highlighting areas where LMMs need further improvement.
@@
Human (Worst) 95.18% 94.24% 96.78% 88.70% 85.49% 82.46% 81.76% 88.91% 92.40%
Human (Best) 95.40% 95.21% 99.42% 95.17% 93.34% 91.73% 84.29% 96.05% 94.02%
@@
MiniCPM-V2.6 (Qwen2-7B) 93.01% 93.22% 93.44% 81.21% 78.31% 77.06% 79.32% 67.86% 84.98%
InternVL2-40B (Nous-Hermes-2-Yi-34B) 94.86% 93.87% 93.56% 80.32% 79.88% 78.01% 77.44% 69.54% 85.17%
Ovis1.5 (Llama3-8B) 92.79% 92.26% 92.12% 80.55% 78.61% 78.59% 78.34% 69.87% 85.08%
LLaVA-OneVision (Qwen2-7B) 92.53% 92.01% 92.07% 81.12% 79.33% 77.98% 79.02% 69.91% 84.88%
CogVLM2-19B (Llama3-8B) 93.31% 92.70% 89.97% 75.41% 64.63% 66.63% 75.88% 61.54% 82.55%
IDEFICS-2 (Mistral-7B-Instruct-v0.2) 89.92% 91.87% 86.50% 75.45% 61.36% 71.04% 73.31% 62.91% 80.14%
DeepSeek-VL-7B 91.48% 91.15% 82.44% 83.73% 63.38% 69.91% 75.40% 60.32% 81.42%
InternLM-XComposer2-VL (InternLM2) 92.79% 95.21% 86.38% 82.64% 68.87% 72.22% 70.77% 64.35% 81.89%
LLaVA-NeXT (Llama3-8B) 92.72% 92.40% 91.15% 83.62% 61.04% 67.07% 76.23% 62.94% 82.88%
LLaVA-NeXT (Qwen-72B) 94.37% 92.72% 91.49% 81.61% 62.40% 73.39% 77.15% 61.44% 83.99%
LLaVA-NeXT (Qwen-110B) 93.83% 91.10% 90.43% 84.71% 67.76% 67.70% 76.25% 64.28% 83.66%
mPLUG-Owl2 (LLaMA-7B) 85.29% 86.26% 83.87% 79.66% 53.73% 57.85% 71.14% 58.47% 76.40%
LLaVA-v1.5 (Vicuna-v1.5-7B) 87.82% 88.65% 83.86% 75.41% 61.39% 65.67% 74.76% 62.69% 78.86%
LLaVA-v1.5 (Vicuna-v1.5-13B) 88.60% 89.57% 86.48% 79.52% 62.33% 58.82% 74.81% 61.56% 79.72%
CogVLM-17B (Vicuna-v1.5-7B) 90.38% 95.17% 85.89% 77.47% 49.56% 47.82% 73.34% 61.34% 78.61%
Qwen-VL (Qwen-7B) 86.14% 86.32% 81.38% 77.47% 52.72% 61.22% 71.61% 57.32% 76.39%
BakLLava (Mistral-7B) 88.91% 81.31% 77.42% 73.81% 52.18% 62.32% 68.37% 49.02% 74.33%
Fuyu-8B (Persimmon-8B) 81.41% 68.27% 66.72% 57.45% 42.24% 48.32% 61.16% 29.65% 63.12%
random guess 32.27% 37.22% 31.03% 42.82% 29.85% 29.78% 26.51% 32.13% 30.80%
@@
The performance results of LMMs on the A-BenchP1 subset, as detailed in Table 1, reveal several key insights: 1) Almost all LMMs show good performance in Basic Recognition, suggesting that they are quite adept at fundamental semantic understanding, which includes recognizing foreground and background objects in AIGIs. 2) However, their effectiveness diminishes in more complex tasks such as Bag-of-Words, particularly in subcategories like Nouns as Adjectives Awareness, Composition Identification, and Number of Objects Counting. These areas require deeper semantic understanding and reasoning, which is critical as users often employ complex prompts that include such nuanced elements. The LMMs’ underperformance here indicates potential challenges in accurately aligning AIGIs with user prompts. 3) Additionally, Outside Knowledge poses significant challenges, with LMMs generally achieving unsatisfactory performance in the Contradiction Overcome subcategory, where AIGIs contain content that defies common sense, requiring LMMs to override their prior knowledge to respond correctly. The subcategory Specific Terms tests the knowledge base of LMMs, where proprietary LMMs generally perform better due to being trained on more recent and extensive datasets. 4) Therefore, to improve the evaluation of text alignment in AIGIs using LMM, it is recommended to simplify overly complex prompts. By employing a divide-and-conquer approach to break down complex prompts into shorter ones, sequential judgment can effectively enhance accuracy.
@@
The performance results of LMMs on the A-BenchP2 subset, as shown in Table 2, illustrate a notable disparity in capabilities: 1) There is a significant performance gap of approximately 23.10% between the top-performing LMMs and human evaluators, highlighting that LMMs lag considerably in quality perception and struggle to accurately assess the quality of AIGIs. 2) Furthermore, most LMMs exhibit their weakest performance in the Generative Distortion Assessment subcategory (except Qwen2-VL-72B), suggesting their ineffectiveness at identifying unexpected generative distortions, such as unnatural appearances and incorrect geometric structures. 3) Interestingly, while humans generally perform better in Technical Quality Perception compared to Aesthetic Quality Evaluation, LMMs show similar performance levels in both subcategories (except Qwen2-VL-72B and MiniCPM-V2.6). This difference likely stems from the more objective nature of technical quality assessments, which leads to more consistent evaluations among humans, whereas aesthetic quality, being more subjective, results in a broader range of opinions and consequently lower performance scores.
@@
Human (Worst) 94.32% 84.49% 86.25% 90.56%
Human (Best) 94.69% 86.01% 93.00% 92.22%